algebraic method to solve?
2.9(x+8.1)=7.8x-3.95

Answer by checkley75(3666)

2.9(x+8.1)=7.8x-3.95
2.9x+23.49=7.8x-3.95
2.9x-7.8x=-3.95-23.49
-4.9x=-27.44
x=-27.44/-4.9
x=5.6 answer.